Greenhab Report Jan 25

Crew 220 GreenHab Report 25-JAN-20

Crew GreenHab Officer: Marufa Bhuiyan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 14.75° C

8:04 A.M.

Floor Unit: 16 °C

12:39 P.M.

Floor Unit: 18 °C

15:10 P.M.

Floor Unit: 8°C

20:30 P.M.

Floor Unit: 17°C

Hours of supplemental light: Light system 7:30 P.M. – 12:00A.M.

Daily water usage for crops: 15.76 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 176.24 Gal.

Time(s) of watering for crops:

8:20 A.M.

6:25 P.M.

SEASONAL PLANTING HISTORY: N/A

Change to crops:

Crops picked for cooking: 2g cilantro, 1g chives, 4g lettuce, 1g lemon balm, 35g icicle radish leaves, 3g acorn pumpkin, 1g rosemary.

OTHER NOTES:

Swept the floor because water was dripping from the shelves and it was stored/clogged in front of the table and blue tank. Next time, I should be more careful when watering

Moved two white bowls with snow pea plants into the floor, as they were drying out probably because of excess heat from the heater

Put some bricks/stones in front of the entrance (below the thermometer just outside of GreenHab), as the entrance was muddy and slippery, there were some risks of falling down

QUESTION: There are some empty tubs with soil, may I use them for harvesting tomorrow? For example: 9 Pepper tubs tags, 2 red tomatoes etc. Please advise. Thanks.

Operations Report Jan 25

Crew 220 Operations Report 25JAN2020

Sol: 6

Name of person filing report: Alyx Skywalker

Non-nominal systems: Primary Generator, Water Heater

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 26 (as reported by Outpost)

From what time last night: 17:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 100%

Diesel Reading – 45%

Station Propane Reading- 86%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0148233,4

Water (static tank) (gal) – 261

Static to loft Pump used – Yes

Water in Green Hab (gal): 184

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Yes, by Outpost

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – No

Starting Hours: 127.9

Beginning Charge: 100%

Ending Hours: 127.9

Ending Charge: 100%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 133.9

Beginning Charge: 100%

Ending Hours: 133.9

Ending Charge: 100%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? Yes. The Habcar was taken to town for water for the GreenHab.

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Changed loft tank water filter. Also took pictures per Outpost’s request of the loft tank’s hoses so they can replace the leaking hose.

Summary of Science Dome operations: Our hab-brewed Kombucha was examined under the microscope.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: Nominal

Questions, concerns, and requests to Mission Support: None.

Sol Summary Report Jan 25

Crew 220 Sol 6 Summary Report 25-01-2020

Sol: 6

Summary Title: Science day in Hab

Author’s name: Johannes Svensoy, MD.

Mission Status: Indoorsily Nominal

Sol Activity Summary: The day started with a joint briefing between the two Crews after MAU Crew came over to the MDRS Hab. Both Crews presented parts of the New Right Stuff book with further discussion around each crew member’s personal mission and how both Crews could support each other here on Mars. After lunch the rest of the day was set for medical training. After doing a trauma exercise with an ankle trauma, the whole Crew was trained in both first and secondary assessment based on the P-MARCH-P protocol. The Crew trained on trauma cases using the training briefed. The whole crew was trained on using the Golden Bubble pressurized medevac device and how to implement this into Medical EVA missions for future medical EVA drills. The Crew has also grown Kombucha in the Science lab, which after a few days of growing was tested and also analyzed for both bacterial growth.

Look Ahead Plan: On Sol 7 there is a planned change of Stations, meaning both Crews will change stations, a few of the crewmembers are going into a simulated quarantine.

Anomalies in work: None

Weather: Overcast, no precipitation. Average 0C, high 2C, low +3C, 10% chance of precipitation.

Crew Physical Status: Well, excited for the days ahead.

EVA: None today.

Reports to be filed:

Sol Summary

Commander’s Report

Photo Report

Operations Report

GreenHab Report

EVA Request

Support Requested: More toilet paper please (down to last rations).

Commander Report Jan 25

MDRS Commander Report 25JAN2020

Sol 6

Author: Shawna Pandya, Commander

Title: A solset on MDRS, a solrise on MAU

Today was a different day – the MDRS and MAU crew rendezvoused at MAU station to focus on crew training, learning and growth. In the morning, the crews worked on personal growth exercises based on the book "The New Right Stuff" and went through the questions in the book. Next, after a team lunch, the medically trained personnel of the team took the rest of the crew through a medical simulation, followed by principles of medical triage in a trauma scenario, and practicing these drills. Next, one of the crew members demonstrated the use of their "Golden Bubble" device, designed to transport injured astronauts in a trauma scenario. The rest of the evening passed by with body-mapping artistic activities, training on a VR device, being briefed on communications in the field and a "Grand Rounds" academic presentation.

Every day on Mars brings its challenges, but plants do not blossom without rain. For this crew’s personal growth, life on Mars -is- the rain.

On a personal note, tomorrow myself and the XO move out to take command of MAU station. We are impressed with what the current crew has achieved, and knowing XO Svensoy and working with him as well as I do, we very much look forward to conducting MAU Operations as a team.

Till the Next Sol,

Commander Pandya

Commander Report Jan 25

MDRS MAU Commander Report 25JAN2020

Sol 6

Title: Learning Medical Trauma Response and Demo New Invention

Author: Commander John Hanacek

Today was the start of crew medical training, learning protocol for first response trauma medicine. We drilled with a scenario of a crew member who hurt their leg. I found my training as a CA state parks beach lifeguard coming back to me, yet I was a little rusty! We are privileged here to have two three medical doctors, two who are experts in trauma medicine, one who works in hospital and one with deep field experience, and an ex special forces soldier to assist in training. I have been looking forward to this mission to grow my skills under the supervision of these professionals in this austere, isolated environment. Many members of the crew have no formal medical training and so this is a fascinating opportunity to test training of life-saving skills. If we are to thrive on Mars, we are required to have intimate knowledge of field medical skills. Plus, one crew member has brought in an invention, “The Golden Bubble,” a pressurized field medical system that is meant to enable the successful stabilization and transport of injured astronauts even if their pressure suits are compromised. We will be testing our skills and this unique technology on future medical EVAs. We are all looking forward to such a unique opportunity.

Tomorrow the MDRS and MAU crews will switch places. This is the next stage of the unique ‘station to station’ experiment going on here and will represent a fresh opportunity to see a new side of the Martian terrain and a new experience for all of us.

Sol Summary Report Jan 25

MDRS MAU Crew 220 Sol Summary Report 25JAN2020

Sol: 6

Summary Title: Sixth Sol Successfully Completed

Author’s name: Morgan Kainu

Mission Status: Healthily Nominal

Sol Activity Summary: MAU crew joined the MDRS crew for medical trauma training at the MDRS hab. We were also briefed on one of the crew member’s science project, the "Golden Bubble".

Look Ahead Plan: Sol 7 will include the switching of stations for the two crews.

Anomalies in work: NA

Weather: Cloudy, low -3 C high 2 C, humidity: 88%, wind: 5 km/h

Crew Physical Status: Nominal

EVA: NA

Reports to be filed: Operations Report, Commander, Photos

Support Requested: None at this stage

Operations Report Jan 25

Crew 220 Operations Report 25-JAN-2020

Sol: 6

Name of person filing report: John Hanacek

Non-nominal systems: N/A

Generator: Not Run

Hours run: 0

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): n/a

Diesel Reading – N/A

Station Propane Reading- 2%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in Green Hab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Diemos rover used: N/A

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – N/A

Starting Hours: N/A

Beginning Charge: N/A

Ending Hours: N/A

Ending Charge: N/A

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: N/A

Starting Hours: N/A

Beginning Charge: N/A

Ending Hours: N/A

Ending Charge: N/A

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: Structural integrity nominal. Floors becoming dirty due to mud. Will be reinforcing with cardboard

Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

EVA Report – January 24th

  

Crew 220 EVA Report 24-01-2020

EVA # 1

Author: Johannes Svensoy, MD

Purpose of EVA: Medical EVA Trauma training

Start time: 0900

End time: 1200

Narrative:

PRE-EVA: The crew woke up eager for our first Medical EVA. The morning before EVA we set up to conduct some science related to crew health in the Science dome. This took more time than anticipated, but all EVA Crew were hydrated and fed before leaving for EVA. But the terrain here on Mars was muddy, as usual.

EVA:

During depressurization one crewmember, had communication problems, this was due to a loose connection and was resolved in the airlock during depressurization. The EVA crew walked over to MAU Station by foot and met up with the MAU EVA Crew. During the MAU briefing one of the crewmembers felt uncomfortable, and it was decided to return said crew member to the hab. Unfortunately, there were ongoing radio and comms problems after this, therefore the remaining crew docked at MDRS Hab for a debrief and testing of communication equipment. The aforementioned crewmember was assessed in Hab and found to feel better.

Destination: MAU Station, Medical EVA site 2 (12s518287,4250596)

Coordinates (use UTM NAD27 CONUS):

Participants: Hab Com: Skywalker. EVA Crew: Johannes, Marufa, Connie

Road(s) and routes per MDRS Map: Entrance road out to Cow Dung Road. Cow Dung Road South to MAU station.

Mode of travel: EVA Walk.

Support Requested: No support requested during EVA.

EVA # 2

Cancelled because of radio communication problems. Decided to rather debrief and go through radio communication protocols.